# Gate tunable enhancement of supercurrent in hybrid planar Josephson junctions

### v1

@@ Abstract
Figure 1: Device geometry and Fraunhofer patterns at different gate configurations. (a) alse-color scanning electron micrograph of the measured device. (b) of the device and the material stacks. (c)(d) Differential resistance as a function of the bias current and out-of-plane magnetic field for (c) JG = 0 V and TG = 0 V and (d) JG = 0 V and TG = -5 V.
@@
We fabricated Josephson junctions on epitaxial superconducting Al thin films grown in-situ on InAs heterostructures. Such devices have shown high transparency [28, 29, 30, 31] a spin-orbit induced anomalous phase [32]. The junction is 4 $\mu$m long with a width of around 100 nm. Two rows of periodic holes are etched on each side of the Al contacts as shown in Fig. 1(a). Each hole is approximately 110 nm wide and 220 nm long. To control the chemical potential in the junction as well as in the hole region, we fabricated two layers of gates. In the first layer, a junction gate (JG) that covers the middle section of the junction is used to control the chemical potential in the junction. In the second layer, which is separated from the first layer of the gate by a second layer of dielectric, a top gate (TG) covers a much larger device region that includes the holes and the ends of the junction. A schematic diagram of the device and the material stacks are presented in Fig. 1(b). As shown in Fig. 1(a), the JG is shorter than the junction by 100 nm at each end by design. While JG itself cannot fully deplete the junction since it does not cover the ends of the junction, the chemical potential in the whole junction can still be fully controlled by using JG and TG together (Fig. S1). We notice TG depletes the ends of the junction around -2.5 V (Fig. S1). Since the 2DEG inside the hole region should have a similar density and coupling to TG as the 2DEG in the junction, we expect the 2DEG in the hole region should also be depleted around -2.5 V by TG. The depletion of the 2DEG in the hole region has a nontrivial effect on the junction, as discussed below. All the measurements in this study are performed in a dilution refrigerator equipped with a three-axis vector magnet. As shown in Fig. 1(a), the z-axis of the magnet is perpendicular to the device plane, while x and y-axes are in-plane fields aligned parallel and perpendicular to the current, respectively. Differential resistance is measured using standard low-frequency lock-in techniques in a four-point manner (more details about the device fabrication and measurement can be found in the methods section).
@@
Figure 2: Supercurrent gate dependence at different in-plane magnetic fields. (a) Differential resistance as a function of the bias current and TG voltages at zero magnetic field. (b) Differential resistance as a function of the bias current and TG voltages at $B_{y}$ = 200 mT. The Supercurrent is significantly enhanced when more negative voltages is applied to TG. (c)(d) Differential resistance as a function of the bias current and JG voltages at $B_{y}$ = 200 mT when TG = 0 V (c) and TG = - 5 V (d). Depletion of holes (d) brings more supercurrent than when the holes are not depleted (c).
@@
Figure 3: Supercurrent in-plane field dependence at different gate configurations. (a) Switching current extracted from panels (b) and (c). (b) Differential resistance as a function of the bias current and $B_{y}$ when JG = 0 V and TG = 0 V. (c) Differential resistance as a function of the bias current and $B_{y}$ when JG = 0 V and TG = -5 V. When the holes are depleted (c), the supercurrent shows an almost linear decreasing with increasing $B_{y}$.
@@
In Fig. 1 (c)(d), we present the differential resistance as a function of the bias current and applied out-of-plane magnetic field for two different gate configurations. To eliminate the hysteresis due to heating effects, the current bias is always swept from zero to high bias in these two scans. All gates are set to 0 for Fig. 1(c). For the results shown in Fig. 1(d), TG is set to -5 V while JG remains at 0. In Fig. 1(d), all the holes are supposed to be depleted as the result of a TG voltage below -2.3 V. In both configurations, the observed Fraunhofer patterns are symmetric with respect to the bias and the out-of-plane field, indicating an absence of hysteresis and a uniform supercurrent distribution across the junction. It is worth noting that extra resistance peaks have been observed in Fig. 1(d), which possibly indicates a stronger multiple Andreev reflection when the hole region is depleted. Overall, the two Fraunhofer patterns show similar periodicity, suggesting the effective junction area is not significantly modified by the depletion of the hole region.
@@
Next, we characterize the supercurrent gate dependence at different magnetic fields. At zero field, the supercurrent decreases when the voltage on TG is swept from 0 to -3 V as the result of the depletion of the ends of the junction (Fig. 2(a)). Below TG = -3 V, however, the supercurrent remains almost unchanged. For JG, the dependence is simpler as the supercurrent always monotonically decreases with decreasing JG voltages (Fig.s). When the magnetic field is turned on, the supercurrent shows a very different behavior. At $B_{y}$ = 200 mT, where $B_{y}$ is perpendicular to the supercurrent, the supercurrent is enhanced when more negative voltage is applied to the TG (Fig. 2(b)). The supercurrent reaches its maximum around TG = -2.5V and remains constant after that. Applying a more negative voltage on TG can deplete both the ends of the junction and the hole region. Depletion of the ends of the junction should lead to a drop of the supercurrent as shown in Fig. 2(a). For fields parallel to the supercurrent, we also observed a similar enhancement of supercurrent with TG voltages(Fig.s ). When TG is fixed to 0 V and -5 V and $B_{y}$ = 200mT, sweeping JG reveals that the supercurrent still monotonically decreases with decreasing JG voltages (Fig. 2(c)(d)). That indicates the enhancement of supercurrent is solely determined by the TG voltage and is not related to the change of chemical potential in the junction.
@@
The enhancement of the supercurrent can also be observed in in-plane magnetic field scans when the gate voltages are fixed. In Fig. 3(b), we plot differential resistance as a function of $B_{y}$ and bias current for TG = 0 V. The supercurrent exhibits a nonlinear behavior as it first quickly decreases at low fields followed by a much slower decline at higher fields. The supercurrent also shows some wiggles at higher fields. Overall, the supercurrent at $B_{y}$ = 0.5 T is less than 10 percent of the supercurrent at $B_{y}$ = 0.1 T. We notice the differential resistance of the ”supercurrent” is finite at high fields, which is occasionally seen in our junctions. We attribute it to some part of the junction becoming normal at high fields. But the sudden dropping in resistance and the Fraunhofer patterns at high fields (Fig. S2 ) indicate there is still supercurrent flow through the semiconductor part of the junction. When TG is fixed to -5 V, Fig. 3(c), the supercurrent has an almost linear dependence on $B_{y}$. From $B_{y}$ = 0.1 T $B_{y}$ Bx = 0.5 T, the supercurrent still preserves 40 percent of its value and the resistance remains to zero. In Fig. 3(a), we extract the switching current from Fig. 3(b)(c) and plot them together. As can be seen, the two switching currents cross around $B_{y}$ = 0.12 T, confirming the observation that the enhancement of the supercurrent only happens at finite fields and when the voltage on TG is below the specific value. In another device, we have observed a similar enhancement of supercurrent at finite fields when TG is below a certain value (see supplementary information for more details).
@@
Figure 5: Profile of $J_{x}$ for the case when $E_{Z}=0$ and $\mu_{h}=\mu$, M1 point in Fig. 4 (c). (b), (c), (d) Profiles of $J_{x}$ for the case when $E_{Z}=1.14\Delta$ and $\mu_{h}$ corresponds to the points N1, N2, N3 in Fig. 4 (d), respectively. $J_{0}$ is the average current density for the case when $E_{Z}=0$, $\mu_{h}=\mu$. The red dashed lines indicate the boundary of the normal strip. The red boxes show positions of some of the depleted holes.
@@
Figure 6: (a) Layout of simplified two-strip model. (b) Scaling of $I_{c}$ with $\mu_{h}$ obtained using the simplified model for the case when $l_{2}=\xi/6$. (c) Same as (b) for the case when $l_{2}=\xi/2$. In (b) and (c) $I_{0}=I_{c}(\mu_{h}=\mu)$.
@@
To understand the origin of this dependence of critical currents on the gate voltage, we setup a tight-binding model for the Boguliobov de Gennes (BdG) Hamiltonian describing the system, see SM for details, using the python package Kwant [33]. To be able to obtain from the model all the desired quantities, in particular the critical current, with the available computational resources, we scaled down all the dimensions while using a value of the superconducting gap $\Delta$ and of the chemical potential $\mu$ in the regions outside the holes such that the ratio between the geometric dimensions and the superconducting coherence length, $\xi$, is of the same order as in the experiment. For the results presented below we use a value of $\Delta$ 16 times larger than the physical one and set $\mu=1.25\Delta$. The geometry of the model used is shown in Fig. 4 (a). To properly model the superconducting leads we chose a value of $L_{x,l}$, see Fig. 4 (a), sufficiently larger than $\xi$ to avoid spurious finite-size effects.
@@
Figure 4 (b) shows the evolution of $I_{c}$ with the strenghth of the Zeeman energy $E_{Z}$ due to an in-plane magnetic field perpendicular to the current for the case when $\mu_{h}=\mu$, in red, and $\mu_{h}=-1.25\Delta$, in green. The Zeeman energy in the superconductor is taken to be $1/2$ the value in the normal and depleted regions. In the first case the holes are normal regions, i.e., regions where $\Delta$ is set to zero, with the same carrier density as the rest of the system. In the second case the carrier density in the holes is lower than in the areas around them. In both cases we see that $I_{c}$ first decreases with $E_{Z}$ up to $E_{Z}\approx 0.6\Delta$. For $E_{z}>0.6\Delta$ in the first case $I_{c}$ changes non-monotonically with $E_{Z}$, a behavior that can be attributed to the almost closing a reopening of the gap of the ABS’s spectrum In the second case $I_{c}$ keeps decreasing also for $E_{z}>0.6\Delta$, albeit more slowly, suggesting that in this case the gap of the ABS doesn’t recover with $E_{Z}$.
As in the experimental case, from Fig. 4 (b), we see that above a threshold value of $E_{Z}$, $E_{Z}\approx 1.1\Delta$, $I_{c}$ for the depleted case is larger than for the non-depleted suggesting that for $E_{Z}>1.1\Delta$ the evolution of $I_{c}$ with $\mu_{h}$ might be not monotonic. This is confirmed by the results for the evolution of $I_{c}$ with respect to $\mu_{h}$ shown in Fig. 4 (c), (d), for the $E_{Z}=0$ and $E_{Z}=1.14\Delta$ cases, respectively. We see that for $E_{Z}=0$ $I_{c}$ decreases monotonically as the depletion ($-\mu_{h}$) of the holes increases, whereas for $E_{Z}=1.14\Delta$ $I_{c}$ varies non-monotonically with $\mu_{h}$, in qualitative agreement with the experimental results.
@@
Figure 5 shows the spatial profile of the quasiparticle current density $J_{x}(x,y)=-i(\hbar/2m^{*})\sum_{\epsilon_{n}<0}(\psi_{n}^{*}\nabla\psi_{n}-\psi_{n}\nabla\psi_{n}^{*})$ [34], where $\psi_{n}$ are the eigenstates of the BdG Hamiltonian. Notice that the regions where $J=0$ are regions where the current is carried by the superconducting condensate. Panel (a) shows the profile of $J$ for the case when $E_{Z}=0$. We see that the presence of the holes induces a periodic modulation in the transverse direction of the current in the normal region of the JJ. Panels (b), (c), (d) show the results for the case $E_{Z}=1.14\Delta$ for the values of $\mu_{h}$ denoted by $N1$, $N2$, and $N3$ in Fig. 4 (d). We see that for the value of $\mu_{h}$ for which $I_{c}$ is maximum, panel (c), $J$ is more uniform in the central region of the JJ. This suggest that the optimal value of $\mu_{h}$ results in a better matching of the quasiparticle wave-functions across the different regions of the JJ.
@@
To check that this is the case we considered a simplified model, shown in Fig. 6 (a). In this model the the holes are effectively replaced by a normal strip, shown in green in Fig. 6 (a), at a distance $l_{2}$ from the normal region of the JJ. Notice that the difference of the superconducting phase across the normal region modeling the holes is set to zero. This is done to take into account that in the experimental geometry there are paths between the holes that connect the different superconducting regions on the same side of the JJ’s normal region, the region shown in orange in Fig. 6 (a), that separates the left and right superconducting leads.
@@
Using the simplified model we were able to see that the critical parameter determining the nature of the evolution of $I_{c}$ with respect to $\mu_{h}$ is the distance $l_{2}$ between the normal region modeling the holes, the depletion strip, and the JJ normal region. Figures 6 (b), (c), show the results for $l_{2}=\xi/6$ and $l_{2}=\xi/2$, respectively, when both $E_{Z}$ and SOC are not present. The results show that when $l_{2}\ll\xi$ $I_{c}$ decreases monotonically with ($-\mu_{h}$), and that when $l_{2}\sim\xi$ $I_{c}$ varies non-monotonically with (-$\mu_{h}$). For $l_{2}\gg\xi$ the two normal strips are effectively decoupled from each other and so $\mu_{h}$ has no effect on $I_{c}$. These results suggest that the main reason why experimentally a non-monotonic scaling of $I_{c}$ vs (-$\mu_{h}$) is observed in the presence of an in-plane magnetic field is the fact that when $E_{Z}\gtrsim\Delta$ the ABSs’ wave functions decay faster in the superconducting regions at the sides of the JJ’s normal region, effect that in the simplified model corresponds to a reduction of $\xi$ [^1] In reality, the magnetic field will also suppress the superconducting gap, however, for the magnetic field considered, such effect is much smaller than the one arising from the reduction of the decay length of the ABS states. and therefore to an increase of the ratio $l_{2}/\xi$ to values $\sim 1$ for which $I_{c}$ scales non-monotonically with $\mu_{h}$.
@@
Project was funded by the US Department of Energy, Office of Basic Energy Sciences, via Award DE-SC0022245. It was partially supported in part by DARPA Topological Excitations in Electronics (TEE) program under grant No. DP18AP900007, the U.S. Office of Naval Research (ONR) through Grants No. N000142112450 and MURI No. N000142212764. H.F. and E.R. thank Joseph J. Cuozzo for very helpful discussions during the initial stages of the work.

### v2

@@ Abstract
Figure 1: Device geometry and Fraunhofer patterns at different gate configurations. (a) False-color scanning electron micrograph of the measured device. (b) schematic of the device and the material stacks. (c),(d) Differential resistance as a function of the bias current and out-of-plane magnetic field for JG = 0 V and TG = 0 V (c) and JG = 0 V and TG = -5 V (d).
@@
We fabricated Josephson junctions on epitaxial superconducting Al thin films grown in-situ on InAs heterostructures. Such devices have shown high transparency [28, 29, 30, 31] and a spin-orbit induced anomalous phase [32]. The Josephson junction is 4 $\mu$m long with a width of around 100 nm. Two rows of periodic holes are etched on each side of the Al contacts together with the junction gap as shown in Fig. 1(a). Each hole is approximately 110 nm wide and 220 nm long. To control the chemical potential in the junction as well as in the hole region, we fabricated two layers of gates. In the first layer, a junction gate (JG) that covers the middle section of the junction is used to control the chemical potential in the junction. In the second layer, which is separated from the first layer of the gate by a second layer of dielectric, a top gate (TG) covers a much larger device region that includes the holes and the ends of the junction. A schematic diagram of the device and the material stacks are presented in Fig. 1(b). As shown in Fig. 1(a), the JG is shorter than the junction by 100 nm at each end by design. While JG itself cannot fully deplete the junction since it does not cover the ends of the junction, the chemical potential in the whole junction can still be fully controlled by using JG and TG together. We notice TG depletes the ends of the junction around -2.5 V (Shown in Fig. S1 in Suppl Materials). Since the 2DEG inside the hole region should have a similar density and coupling to TG as the 2DEG in the junction, we expect that the 2DEG in the hole region should also be depleted around -2.5 V by TG. Tuning the chemical potential of the 2DEG in the hole region has a nontrivial effect on the junction, as discussed below. All the measurements in this study are performed in a dilution refrigerator equipped with a three-axis vector magnet. As shown in Fig. 1(a), the z-axis of the magnet is perpendicular to the device plane, while x and y-axes are in-plane fields aligned parallel and perpendicular to the current, respectively. Differential resistance is measured using standard low-frequency lock-in techniques in a four-point manner (more details about the device fabrication and measurement can be found in the methods section).
@@
Figure 2: Supercurrent gate dependence at different in-plane magnetic fields. (a) Differential resistance as a function of the bias current and TG voltages at zero magnetic field. (b) Differential resistance as a function of the bias current and TG voltages at $B_{y}$ = 200 mT. The Supercurrent is significantly enhanced when a more negative voltage is applied to TG. (c),(d) Differential resistance as a function of the bias current and JG voltages at $B_{y}$ = 200 mT when TG = 0 V (c) and TG = - 5 V (d). Supercurrent is always monotonically decreasing with decreasing JG voltages.
@@
Figure 3: Supercurrent in-plane field dependence at different gate configurations. (a) Switching current extracted from panels (b) and (c). (b) Differential resistance as a function of the bias current and $B_{y}$ when JG = 0 V and TG = 0 V. (c) Differential resistance as a function of the bias current and $B_{y}$ when JG = 0 V and TG = -5 V, the supercurrent shows an almost linear decreasing with increasing $B_{y}$.
@@
In Fig. 1 (c)(d), we present the differential resistance as a function of the bias current and applied out-of-plane magnetic field for two different gate configurations. To eliminate the hysteresis due to heating effects, the current bias is swept from zero to high bias in these two scans. All gates are set to 0 for Fig. 1(c). For the results shown in Fig. 1(d), TG is set to -5 V while JG remains at 0. In Fig. 1(d), all the holes are supposed to be depleted as the result of a TG voltage below -2.5 V. In both configurations, the observed Fraunhofer patterns are symmetric with respect to the bias and the out-of-plane field, indicating an absence of hysteresis and a uniform supercurrent distribution across the junction. It is worth noting that extra resistance peaks have been observed in Fig. 1(d), which possibly indicates a stronger multiple Andreev reflection when the hole region is depleted. Overall, the two Fraunhofer patterns show similar periodicity, suggesting the effective junction area is not significantly modified by the depletion of the hole region.
@@
Next, we characterize the supercurrent gate dependence at different magnetic fields. At zero field, the supercurrent decreases when the voltage on TG is swept from 0 to -2.5 V as the result of the depletion of the ends of the junction (Fig. 2(a)). Below TG = -2.5 V, however, the supercurrent remains almost unchanged. This is consistent with the result in Fig.s1 which shows that TG depletes the ends of the junction around -2.5 V. For JG, the dependence is simpler as the supercurrent always monotonically decreases with decreasing JG voltages (Fig.s1(c)(d)). When the magnetic field is turned on, the supercurrent shows a very different behavior regarding TG voltages. At $B_{y}$ = 200 mT, where $B_{y}$ is perpendicular to the supercurrent, the supercurrent is significantly enhanced when a more negative voltage is applied to the TG (Fig. 2(b)). The supercurrent reaches its maximum around TG = -2.5V and remains almost constant for TG voltages below that value. As the depletion of the ends of the junction around TG = -2.5 V should only lead to a decrease of the supercurrent instead of an increase of the supercurrent, we suspect the abnormal enhancement of the supercurrent at finite fields could be related to the decreased chemical potential in the TG-controlled hole region. For fields parallel to the supercurrent, we observed a similar enhancement of supercurrent with TG voltages (Fig.s1(b)). When TG is fixed to 0 V and -5 V and $B_{y}$ = 200 mT, sweeping JG reveals that the supercurrent still monotonically decreases with decreasing JG voltages (Fig. 2(c)(d)). That indicates the enhancement of supercurrent is solely determined by the change of the chemical potential in the hole region and is not related to the change of the chemical potential in the junction.
@@
The enhancement of the supercurrent can also be observed in in-plane magnetic field scans when the gate voltages are fixed. In Fig. 3(b), we plot differential resistance as a function of $B_{y}$ and bias current for TG = 0 V and JG = 0 V. The supercurrent exhibits a nonlinear behavior as it first quickly decreases at low fields followed by a much slower decline at higher fields. The supercurrent also shows some wiggles at higher fields. Overall, the supercurrent at $B_{y}$ = 0.5 T is less than 10 percent of the supercurrent at $B_{y}$ = 0.1 T. We notice the differential resistance of the ”supercurrent” is finite at high fields, which is occasionally seen in our junctions. We attribute it to some part of the junction becoming normal at high fields. But the sharp drop in resistance and the Fraunhofer patterns at high fields (Fig.s2 ) indicate there is still supercurrent flow through the semiconductor part of the junction. When TG is fixed to -5 V, Fig. 3(c), the supercurrent has an almost linear dependence on $B_{y}$. From $B_{y}$ = 0.1 T to $B_{y}$ = 0.5 T, the supercurrent still preserves 40 percent of its value and the resistance remains to zero. In Fig. 3(a), we extract the switching current from Fig. 3(b)(c) and plot them together. As can be seen, the two switching currents cross around $B_{y}$ = 0.12 T, confirming the observation that the enhancement of the supercurrent only happens at finite fields and when the voltage on TG is below the specific value. In another device, we have observed a similar enhancement of supercurrent at finite fields when TG is below a certain value (see Supplemental Material (SM) for more details).
@@
Figure 5: (a) Profile of $J_{x}$ for the case when $E_{Z}=0$ and $\mu_{h}=\mu$, M1 point in Fig. 4 (c). (b), (c), (d) Profiles of $J_{x}$ for the case when $E_{Z}=1.14\Delta$ and $\mu_{h}$ corresponds to the points N1, N2, N3 in Fig. 4 (d), respectively. $J_{0}$ is the average current density for the case when $E_{Z}=0$, $\mu_{h}=\mu$. The red dashed lines indicate the boundary of the normal strip. The red boxes show positions of some of the depleted holes.
@@
Figure 6: (a) Layout of simplified two-strip model. (b) Scaling of $I_{c}$ with $\mu_{h}$ obtained using the simplified model for the case when $l_{2}=\xi/2$. $I_{0}=I_{c}(\mu_{h}=\mu)$.
@@
To understand the origin of this dependence of critical currents on the gate voltage, we setup a tight-binding model for the Boguliobov de Gennes (BdG) Hamiltonian describing the system (see SM for details), using the python package Kwant [33]. To be able to obtain from the model all the desired quantities, in particular the critical current, with the available computational resources, we scaled down all the dimensions while using a value of the superconducting gap $\Delta$ and of the chemical potential $\mu$ in the regions outside the holes such that the ratio between the geometric dimensions and the superconducting coherence length, $\xi$, is of the same order as in the experiment. For the results presented below we use a value of $\Delta$ 16 times larger than the physical one and set $\mu=1.25\Delta$. The geometry of the model used is shown in Fig. 4 (a). To properly model the superconducting leads we chose a value of $L_{x,l}$, see Fig. 4 (a), sufficiently larger than $\xi$ to avoid spurious finite-size effects.
@@
Figure 4 (b) shows the evolution of $I_{c}$ with the strenghth of the Zeeman energy $E_{Z}$ due to an in-plane magnetic field perpendicular to the current for the case when $\mu_{h}=\mu$, in red, and $\mu_{h}=-1.25\Delta$, in green. The Zeeman energy in the superconductor is taken to be $1/2$ the value in the normal and depleted regions. In the first case the holes are normal regions, i.e., regions where $\Delta$ is set to zero, with the same carrier density as the rest of the system. In the second case the carrier density in the holes is lower than in the areas around them. In both cases we see that $I_{c}$ first decreases with $E_{Z}$ up to $E_{Z}\approx 0.6\Delta$. For $E_{z}>0.6\Delta$ in the first case $I_{c}$ changes non-monotonically with $E_{Z}$, a behavior that can be attributed to the almost closing and reopening of the gap of the ABS’s spectrum In the second case $I_{c}$ keeps decreasing also for $E_{z}>0.6\Delta$, albeit more slowly, suggesting that in this case the gap of the ABS doesn’t recover with $E_{Z}$.
As in the experimental case, from Fig. 4 (b), we see that above a threshold value of $E_{Z}$, $E_{Z}\approx 1.1\Delta$, $I_{c}$ for the depleted case is larger than for the non-depleted suggesting that for $E_{Z}>1.1\Delta$ the evolution of $I_{c}$ with $\mu_{h}$ might be not monotonic. This is confirmed by the results for the evolution of $I_{c}$ with respect to $\mu_{h}$ shown in Fig. 4 (c), (d), for the $E_{Z}=0$ and $E_{Z}=1.14\Delta$ cases, respectively. We see that for $E_{Z}=0$, $I_{c}$ decreases monotonically as the depletion ($-\mu_{h}$) of the holes increases, whereas for $E_{Z}=1.14\Delta$, $I_{c}$ varies non-monotonically with $\mu_{h}$, in qualitative agreement with the experimental results.
@@
Figure 5 shows the spatial profile of the quasiparticle current density $J_{x}(x,y)=-i(\hbar/2m^{*})\sum_{\epsilon_{n}<0}(\psi_{n}^{*}\nabla\psi_{n}-\psi_{n}\nabla\psi_{n}^{*})$ [34], where $\psi_{n}$ are the eigenstates of the BdG Hamiltonian and $m^{*}$ is the effective mass. Notice that the regions where $J=0$ are regions where the current is carried by the superconducting condensate. Panel (a) shows the profile of $J$ for the case when $E_{Z}=0$. We see that the presence of the holes induces a periodic modulation in the transverse direction of the current in the normal region of the JJ. Panels (b), (c), (d) show the results for the case $E_{Z}=1.14\Delta$ for the values of $\mu_{h}$ denoted by $N1$, $N2$, and $N3$ in Fig. 4 (d). We see that for the value of $\mu_{h}$ for which $I_{c}$ is maximum, panel (c), $J$ is more uniform in the central region of the JJ. This suggest that the optimal value of $\mu_{h}$ results in a better matching of the quasiparticle wave-functions across the different regions of the JJ.
@@
To check that this is the case we considered a simplified model, shown in Fig. 6 (a). In this model the holes are effectively replaced by a normal strip, shown in green in Fig. 6 (a), at a distance $l_{2}$ from the normal region of the JJ. Notice that the difference of the superconducting phase across the normal region modeling the holes is set to zero. This is done to take into account that in the experimental geometry there are paths between the holes that connect the different superconducting regions on the same side of the JJ’s normal region, the region shown in orange in Fig. 6 (a), that separates the left and right superconducting leads.
@@
Using the simplified model we were able to see that the critical parameter determining the nature of the evolution of $I_{c}$ with respect to $\mu_{h}$ is the distance $l_{2}$ between the normal region modeling the holes, the depletion strip, and the JJ normal region. Figure 6(b) shows the result for $l_{2}=\xi/2$, when both $E_{Z}$ and SOC are not present. We find that when $l_{2}\ll\xi$, $I_{c}$ decreases monotonically with ($-\mu_{h}$), while at $l_{2}\sim\xi$, $I_{c}$ varies non-monotonically with (-$\mu_{h}$) as shown in Fig. 6(b). For $l_{2}\gg\xi$ the two normal strips are effectively decoupled from each other and so $\mu_{h}$ has no effect on $I_{c}$. These results suggest that the main reason why experimentally a non-monotonic scaling of $I_{c}$ vs (-$\mu_{h}$) is observed in the presence of an in-plane magnetic field is the fact that when $E_{Z}\gtrsim\Delta$ the ABSs’ wave functions decay faster in the superconducting regions at the sides of the JJ’s normal region, effect that in the simplified model corresponds to a reduction of $\xi$ [^1] In reality, the magnetic field will also suppress the superconducting gap, however, for the magnetic field considered, such effect is much smaller than the one arising from the reduction of the decay length of the ABS states. and therefore to an increase of the ratio $l_{2}/\xi$ to values $\sim 1$ for which $I_{c}$ scales non-monotonically with $\mu_{h}$.
@@
Project was funded by the US Department of Energy, Office of Basic Energy Sciences, via Award DE-SC0022245. H.F. and E.R. thank Joseph J. Cuozzo for very helpful discussions during the initial stages of the work.